A set of scores ranges from a high of x=24 to a low of x=12 and has a mean of 18. Which of the following is the most likely value for the standard deviation for these scores? Please show work. (The answer is 3 points, but I don't understand why.)

a. 3 points

b. 6 points

c. 12 points

d. 24 points
Answer #1

Using the thumb rule for calculating the standard deviation

std. dev. = (max - min)/4

std. dev. = (24 - 12)/4 = 3

Hence, answer is 3 points